OpenStudy (anonymous):

For Distance measurement, one gigameter GM equals to 10^9 m. If one pentametre(pm) is 6 orders of magnitude longer than the gigametre, what is the value of the pentameter in meter?

OpenStudy (anonymous):

i guess 10^12m

OpenStudy (anonymous):

oh it is penta umm maybe 10^5

OpenStudy (anonymous):

i got 10^15 but i had to convert to pentameter so......

OpenStudy (anonymous):

oh lol i didn't read the question completely sorry i messed it up i didn't read about the condition

OpenStudy (anonymous):

no worries

OpenStudy (anonymous):

yup 9+6=15

OpenStudy (anonymous):

i guess you are right

OpenStudy (anonymous):

but i had to convert to gigameter to pentameter in the end

OpenStudy (anonymous):

then pentameter to meter

OpenStudy (anonymous):

10^15 pentameter to meters would be?

OpenStudy (kira_yamato):

1Pm = 10^(9+6) m = 10^15 m

OpenStudy (anonymous):

^^^